# On Signal Peak Power Constraint of Over-the-Air Federated Learning

### v1

@@ III-A Peak-to-Average Power Ratio
In practice, only digital discrete samples can be employed to estimate PAPR. It has been shown that samples sampled at Nyquist rate inaccurately captures PAPR.
Oversampling: To accurately capture PAPR in the time-domain OFDM signal, oversampling is essential. An oversampling factor $L_{\mathrm{os}}$ (typically $L_{\mathrm{os}}=4$ is sufficient [21]) is applied by zero-padding the frequency-domain signal before the IDFT. The oversampled time-domain signal is generated as:
$s_{k,\ell}^{(\mathrm{os})}[i]=\!\!\!\sum_{m=0}^{L_{\mathrm{os}}M-1}\!\!G_{k,\ell}[m]e^{j\frac{2\pi mi}{L_{\mathrm{os}}M}},\quad i\!=\!0,\ldots,L_{\mathrm{os}}M\!-\!1,\vskip-2.84544pt$ (23)
where the zero-padded frequency-domain vector is defined as:
$\hskip-7.11317ptG_{k,\ell}[m]\!=\!\begin{cases}g_{k}[\ell M\!+\!m]b_{k,\ell M+m}^{(in)},&0\!\leq\!m\!<\!M,\\ 0,&M\!\leq\!m\!<L_{\mathrm{os}}M.\end{cases}\vskip-2.84544pt$ (24)
The per-symbol PAPR can therefore be estimated as
@@ III-B Clipping with Oversampling
The commonly used technique to deal with the high peak-to-average power ratio issue is applying the procedure of iterative clipping-and-filtering. Since clipping results in in-band and out-of-band distortions in multi-carrier OFDM [19, 20]. A low-pass filter is applied afterwards to limit out-of-band distortion, which may result in the rise of PAPR, necessitating repeated such procedures.
@@ III-B2 Multi-Carrier OFDM
  • Out-of-band emissions: spectral regrowth on the zero-padded subcarriers $M\leq m<L_{\mathrm{os}}M$ due to clipping, which causes interference to adjacent channels.
@@
To suppress out-of-band emissions, a low-pass filteris applied. Define a rectangular filter:
$H_{\mathrm{filter}}[m]=\begin{cases}1,&0\leq m<M,\\ 0,&M\leq m<L_{\mathrm{os}}M.\end{cases}\vskip-2.84544pt$ (29)
The filtered frequency-domain signal is:
$X_{k,\ell}^{(\mathrm{filt})}[m]=X_{k,\ell}[m]\cdot H_{\mathrm{filter}}[m].\vskip-2.84544pt$ (30)
Converting back to the time domain yields the filtered signal:
$x_{k,\ell}^{(\mathrm{filt})}[i]=\mathrm{IDFT}\bigl\{X_{k,\ell}^{(\mathrm{filt})}[m]\bigr\}[i].\vskip-4.26773pt$ (31)
However, filtering may reintroduce peak regrowth, as it restores some of the constructive interference among subcarriers that clipping had disrupted [19].
@@
To jointly control PAPR and out-of-band emissions, an iterative clipping-and-filtering procedure is commonly employed [22]. The algorithm proceeds as follows for each symbol $\ell$:
@@
    While $x_{k,\ell}^{(j,\mathrm{clip})}[i]$ has out-of-band distortion higher than a threshold:
    1. (a)
      Clip: $x_{k,\ell}^{(j,\mathrm{clip})}[i]=\mathcal{C}_{A_{\max}}\bigl(x_{k,\ell}^{(j-1)}[i]\bigr)$.
    2. (b)
      DFT: $X_{k,\ell}^{(j)}[m]=\mathrm{DFT}\bigl\{x_{k,\ell}^{(j,\mathrm{clip})}[i]\bigr\}[m]$.
    3. (c)
      Filter: $X_{k,\ell}^{(j,\mathrm{filt})}[m]=X_{k,\ell}^{(j)}[m]\cdot H_{\mathrm{filter}}[m]$.
    4. (d)
      IDFT: $x_{k,\ell}^{(j)}[i]=\mathrm{IDFT}\bigl\{X_{k,\ell}^{(j,\mathrm{filt})}[m]\bigr\}[i]$.
@@
    Output: $x_{k,\ell}^{(\mathrm{ICF})}[i]=x_{k,\ell}^{(J_{\max})}[i]$.
The final transmitted signal is $x_{k,\ell}^{(\mathrm{ICF})}[i]$, which satisfies the peak-power constraint with a certain threshold for out-of-band distortion. The flow diagram of the whole procedure can be found in Fig. 2.
@@ IV-A Simulation Settings
[…]
@@ IV-B Simulation Results and Discussions
Peak-to-average power ratio is shown in Fig. 3. Without clipping, the PAPR is high (about $35\text{\,}\mathrm{dB}$), while after clipping and filtering, the PAPR remains close to $15\text{\,}\mathrm{dB}$. Note that the results without clipping and before clipping are different since PAPR depends on the actual values to be transmitted and clipping and filtering procedure changes the actual values to approximate the average of gradient (5). We also observe that single-carrier transmission can exhibit higher PAPR than multi-carrier OFDM transmission, which is unusual for digital modulation schemes. This stems from the nature of amplitude modulation (for the single-carrier case), where the power is directly related to the actual values to be transmitted, which can actually go up to “infinity” if the gradient values can take arbitrarily large values. The decrease in PAPR in the single-carrier case w.r.t. communication rounds may be due to that the model has been well trained, so the gradient values divergence is smaller.
@@
[…]
@@ V Conclusion
This work highlights an overlooked practical issue in over-the-air computation-federated learning: the instantaneous peak power constraint imposed by the nonlinearity in power amplifiers. This is an issue in over-the-air computation-federated learning since it uses analog transmission and some works consider multi-carrier OFDM transmissions. We applied the classic PAPR reduction methods of iterative amplitude clipping and frequency-domain filtering. Enforcing peak power constraint can indeed degrade aggregation and final accuracy, in some cases, the effects are marginal, and in some other cases (multi-carrier with low noise level), cause divergence. Counterintuitively, the impact is stronger at lower noise for multi-carrier OFDM because the scaling factor from optimization is higher and therefore the in-band distortion is too much amplified. This work should motivate further research in peak power constraint-aware power control and better PAPR reduction techniques for single-carrier and multi-carrier over-the-air computation-federated learning systems.

### v2

@@ III-A Peak-to-Average Power Ratio
In practice, Nyquist-rate samples underestimate the true PAPR. An oversampling factor $L_{\mathrm{os}}=4$ [21] is applied by zero-padding the frequency-domain signal ($G_{k,\ell}[m]=g_{k}[\ell M+m]b_{k,\ell M+m}^{(in)}$ for $0\leq m<M$, zero otherwise) before the IDFT:
$s_{k,\ell}^{(\mathrm{os})}[i]=\!\!\!\sum_{m=0}^{L_{\mathrm{os}}M-1}\!\!G_{k,\ell}[m]e^{j\frac{2\pi mi}{L_{\mathrm{os}}M}},\quad i\!=\!0,\ldots,L_{\mathrm{os}}M\!-\!1.\vskip-2.84544pt$ (18)
The per-symbol PAPR is then estimated as
@@ III-B Clipping with Oversampling
The most classic technique without changing the modulation schemes (that are necessary for over-the-air computation) to deal with the high peak-to-average power ratio issue is applying the procedure of iterative clipping-and-filtering [19]. Since clipping results in in-band and out-of-band distortions in multi-carrier OFDM [20], a low-pass filter is applied afterwards to limit out-of-band distortion, which may result in the rise of PAPR, necessitating repeating such procedures.
@@ III-B2 Multi-Carrier OFDM
  • Out-of-band emissions: spectral regrowth on the zero-padded subcarriers $M\leq m<L_{\mathrm{os}}M$ due to clipping, which is to be strictly limited since it may cause interference to adjacent frequency bands.
@@
To suppress out-of-band emissions, a rectangular low-pass filter ($H_{\mathrm{filter}}[m]=1$ for $0\leq m<M$, else 0) is applied in the frequency domain and converted back via IDFT:
$x_{k,\ell}^{(\mathrm{filt})}[i]=\mathrm{IDFT}\bigl\{X_{k,\ell}[m]\cdot H_{\mathrm{filter}}[m]\bigr\}[i].\vskip-2.84544pt$ (23)
However, filtering may reintroduce peak regrowth [19].
@@
To jointly control PAPR and out-of-band emissions, an iterative clipping-and-filtering procedure is commonly employed [22]:
@@
    Repeat until out-of-band distortion is below threshold: clip $x^{(j-1)}$ to $A_{\max}$; apply DFT; apply $H_{\mathrm{filter}}$; apply IDFT to obtain $x^{(j)}$.
@@
    Output $x_{k,\ell}^{(\mathrm{ICF})}[i]=x_{k,\ell}^{(J_{\max})}[i]$.
The flow diagram is shown in Fig. 2.
@@ IV-A Simulation Settings
In the simulations the LeNet convolutional neural network with $62006$ trainable parameters is trained on the Cifar-10 dataset for $500$ communication rounds, where the data is independent and identically distributed among $K=40$ user equipments. In each communication round all user equipments participate and are uniformly distributed on a $100$ meter radius disk around the base station. As a large-scale fading model, the free-space-path loss is used, which is combined with Rayleigh fading as a small-scale fading model. The constraint of the maximum average power $P_{\mathrm{avg},\max}$ of user equipments is chosen to be $23\text{\,}\mathrm{dBm}$, while the constraint of the the maximum instantaneous power $P_{\mathrm{inst},\max}$ is chosen to be $26\text{\,}\mathrm{dBm}$. In multi-carrier OFDM transmission, $M=32$ subcarriers are used in order to reduce OFDM complexity, where each subcarrier transmits with a bandwidth of $60\text{\,}\mathrm{kHz}$. Clipping and filtering iterate until out-of-band emissions are below $10\text{\,}\mathrm{dBm}$. The learning rate is set at 1 and the local batch size at 256.
@@ IV-B Simulation Results and Discussions
Peak-to-average power ratio is shown in Fig. 3. Without clipping, the PAPR is high (about $35\text{\,}\mathrm{dB}$), while after clipping and filtering, the PAPR remains close to $15\text{\,}\mathrm{dB}$. Note that the results without clipping and before clipping are different since PAPR depends on the actual values to be transmitted and clipping and filtering procedure changes the actual values to approximate the average of gradient (2). We also observe that single-carrier transmission can exhibit higher PAPR than multi-carrier OFDM transmission, which is unusual for digital modulation schemes. This stems from the nature of amplitude modulation (for the single-carrier case), where the power is directly related to the actual values to be transmitted, which can actually go up to “infinity” if the gradient values can take arbitrarily large values. The decrease in PAPR in the single-carrier case w.r.t. communication rounds may be due to that the model has been well trained, so the gradient value variance is smaller.
@@
The results on FL test accuracy are shown in Fig. 5. For single-carrier (Fig. 5a), we observe consistent but small gaps between clipped and unclipped cases, with higher noise yielding lower accuracy overall. A noticeable accuracy drop appears at around 50 communication rounds for the clipped case. For multi-carrier OFDM (Fig. 5b), as predicted by the TSE, the clipped case performs worse at lower noise; at $\sigma_{n}^{2}/\mathrm{Hz}=\,$$110\text{\,}\mathrm{dBm}$/Hz, training may even diverge. The corresponding accuracy loss due to clipping (considering the peak power constraint) is shown in Fig. 6. At the later stage of training, the loss of test accuracy is about 1.5–2%, while around 50 rounds it can peak at up to 8% for the drop of accuracy mentioned previously.
@@ V Conclusion
This work highlights an overlooked practical issue in over-the-air computation-federated learning: the instantaneous peak power constraint imposed by the nonlinearity in power amplifiers. This is a practical issue in over-the-air computation-federated learning since amplitude-modulated analog transmission is by itself a potentially high peak-to-average power ratio signal (depending on the distributions of values to be transmitted) and multi-carrier over-the-air computation-OFDM would inherit the high PAPR issue of OFDM. As the first work that reveals this problem in AirComp-FL, in order to confirm that this is actually a problem to be solved, we applied the most classic PAPR reduction method of iterative amplitude clipping and frequency-domain filtering that constrains the signal within its PAPR limits without causing distortions to adjacent frequency bands. We observe that enforcing peak power constraint can indeed degrade aggregation and final accuracy, in some cases, the effects are marginal, and in some other cases (multi-carrier with low noise level), cause divergence. Counterintuitively, the impact is stronger at lower noise for multi-carrier OFDM because the scaling factor from optimization is higher and therefore the in-band distortion is too much amplified. This work should motivate further research in peak power constraint-aware power control and better PAPR reduction techniques for single-carrier and multi-carrier over-the-air computation-federated learning systems.